In a college, out of 4320 students, 2300 are girls. Find the ratio of
 A number of boys to the number of girls.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
Total number of students = 4320
Number of girls = 2300
Number of boys = 4320–2300
= 2020
Ratio of number of boys to the number of girls = 2020 / 2300
= (2 × 2 × 5 × 101) / (2 × 2 × 5 × 115)
= 101 / 115

Thanks